# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ 3.1 Coefficient functions from IR renormalized bubble chain
Clearly, all contributions are absolutely convergent. The factor $\frac{gF}{1+gF}$ clearly has the interpretation of a massive chain of bubbles starting with bubble number $n=1$. The $n=0$ contribution is logarithmically divergent and subtracted by $\delta m^{2}$. We would like to examine the OPE expansion of the above. First notice that the last term in Eq. (89) is due to the $\frac{1}{N}$ order tadpole contribution and contains no factorial behavior at all, so we focus on the the first term.
By investigation the regions, the operators beside the identify operator that we will need are the following:
@@
    Tree-level operators $\prod_{l}\frac{1}{N}\Phi^{a}(i\partial)^{2k_{l}}\Phi^{a}$. At the order $\frac{1}{N}$, these operators are responsible for the IR sub-contributions within the bubble chain when $k^{2}$ itself is in the hard region, as well as the IR contribution on the $(p-k)$ line when $p-k$ becomes soft (this means $k$ is hard). Clearly, these “tree-level” insertions are mutually independent and can be performed separately for the individual bubbles. At the external legs, there can be an arbitrary numbers of tree-level $\frac{1}{N}\Phi^{a}\Phi^{a}$ insertions that always combine together with the $m_{r}^{2}$ insertion. These tree-level contributions can all be removed by simply setting $\mu^{2}=m^{2}$.
@@
    The non-trivial operators $O_{n}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n-2}\Phi^{2}$ with $n\geq 1$. Clearly, these operators are responsible for the contributions when $k$ becomes soft. On the other hand, their coefficient functions are relatively simple and can be obtained by Taylor expanding the $\frac{1}{(p-k)^{2}+m^{2}}$ at small $k$. For example, without mass insertion one has in the symmetric state after angular average of $k$:
@@
    and with mass insertions one can generalize without difficulty. These coefficient functions are clearly absent from any factorial growth.
To conclude, to see the factorial growth, it is sufficient to focus on the coefficient function for the identity operator. We can write the $\frac{1}{N}$ order coefficient function of the identity operator for $S(p^{2})$ at the the power $\frac{1}{(p^{2})^{n+2}}$ in the general form as
$C^{(n)}_{\rm I,NL}(p^{2})=\frac{g^{n+1}}{(p^{2})^{n+2}}\sum_{k=0}^{n+1}\left(\frac{m_{r}^{2}}{g}\right)^{k}C_{n,k}(l_{\mu})\ ,$ (93)
@@ 3.1.1 Calculation of $C_{n,0}$: factorial growth from $\frac{1}{\epsilon}\times\epsilon$
Given the above, let’s consider the calculation detail. For individual bubble chain, one needs the contribution in Eq. (41)
$\frac{1}{2}C_{1}=\frac{1}{p^{2}}\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}f(\epsilon)\times\frac{-1}{4\pi\epsilon}\ ,$ (94)
@@
Given the above, the bare coefficient function (in the below we use hat to denote bare or un-subtracted coefficient functions) corresponding to the $n$-bubble chain reads
@@
The crucial property of the above is that it is analytic at $s=0$ and $\epsilon=0$. Now, the IR subtraction for the sub-diagram can be performed by replacing $s\rightarrow(n+1-k)\epsilon$ with $k$ ranging from 00 to $n$, thus the partly renormalized diagram reads
$R_{b}C_{n,0}=-\frac{1}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}\ .$ (99)
Clearly, here we simply add back the $\frac{1}{4\pi\epsilon}$ counter-terms for $\frac{1}{N}\Phi^{2}$ insertion replacing one of two lines in the bubbles. Now, using the standard summation formulas Palanques-Mestre and Pascual (1984); Beneke and Braun (1994, 1995), as well as the new one
@@
one has up to all constant terms,
$R_{b}C_{n,0}$
@@
This term must be canceled by subtracting the IR contribution from the $p-k$ propagator. For that one needs
$R_{p-k}\tilde{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\frac{H(\epsilon,s)}{\epsilon^{n+1}}\ ,$ (104)
@@
The IR subtraction leads to
$R_{b}R_{p-k}\tilde{C}_{0}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{1}{\epsilon}\frac{d^{n}}{ds^{n}}H(\epsilon,s)\bigg{|}_{s=0}+\frac{n+2}{2}\frac{d^{n+1}H(\epsilon,s)}{ds^{n+1}}\bigg{|}_{s=0}\bigg{)}.$ (106)
For this, one needs the
$H(\epsilon,s)=\left(\frac{\mu^{2}}{p^{2}}\right)^{s}-\frac{\epsilon}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\left(12\ln\left(\frac{\mu^{2}}{p^{2}}\right)+\pi^{2}s\right)\ .$ (107)
@@
which indeed cancel with that for $R_{b}C_{0}$. Now, let’s consider the $p$-independent constant terms that has to cancel with the operator renormalization of $O_{n}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n-2}\Phi^{2}$. One writes the constant part as
$-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{(-1)^{n}}{(n+1)\epsilon^{n+2}}G(n,\epsilon,0)+Z\left(\frac{1}{\epsilon}\right)\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{)}\ .$ (109)
Now, for $n\in Z_{\geq 1}$ one can rewrite the above as
$-\frac{1}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{(}Z+\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ .$ (110)
@@
$Z=-\text{poles of}\ \bigg{(}\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ ,$ (111)
@@
Clearly, this part is free from any factorial growth since the bracket is analytic in the neighborhood of $\epsilon=0$, thus the Taylor expansion coefficients can only growth exponentially.
@@
Notice that the first two terms arise due to $\frac{1}{\epsilon}\times\epsilon$ effects. The total result reads
$C_{n,0}=C_{n,0}^{A}+C_{n,0}^{B}\ .$ (118)
For $n=1$, one obtains $C_{1,0}^{A}=0$, while
@@
$C_{n,0}^{\rm asy1}=-\frac{1}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{\mu^{2}}{p^{2}}\bigg{)}\times\left(1+\frac{\mu^{2}}{2^{n+1}p^{2}}+..\right)\ .$ (122)
It is interesting to notice that the leading factorial growth is again proportional to a high-power correction. Here there are two possibilities. The first possibility is that they cancel within the same coefficient function for the identity operator but with mass-insertions. The second possibility is that they cancel with the operator contribution for $O_{n+1}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n}\Phi^{2}$ and $O_{n+2}$…., but with the same bubble-chain number, because this diagram contains power-divergences.
@@
$C_{n,0}^{\rm asy2}=\frac{(-1)^{n}}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{p^{2}}{\mu^{2}}\bigg{)}\times\left(1+\frac{p^{2}}{2^{n+1}\mu^{2}}+...\right)\ .$ (123)
They will cancel with factorial growths for operators $O_{n-1},O_{n-2}...$ but with the same numbers of bubbles.

### v2

@@ 3.1 Coefficient functions from IR renormalized bubble chain
Clearly, all contributions are absolutely convergent. The factor $\frac{gF}{1+gF}$ clearly has the interpretation of a massive chain of bubbles starting with bubble number $n=1$. The $n=0$ contribution is logarithmically divergent and subtracted by $\delta m^{2}$. We would like to examine the OPE expansion of the above. First notice that the last term in Eq. (89) is due to the $\frac{1}{N}$ order tadpole contribution and contains no factorial behavior at all, so we focus on the the first term. See Fig. 5 for a depiction of the massive bubble chain diagram with $n=5$.
Figure 5: The massive bubble-chain diagram with $n=5$ bubbles. Solid lines are massive scalar propagators. Notice in the figure we have separated the quartic interaction vertices slightly apart just to show the contractions among the $O(N)$ indices, no point-spiting-type regularization has been implied. At the order $\frac{1}{N}$ for the two-point function $S(p^{2})$, the $O(N)$ indices only contract within each bubbles.
By investigating the regions, the operators beside the identify operator that we will need are the following:
@@
    Tree-level operators $\prod_{l}\frac{1}{N}\Phi(i\partial)^{2k_{l}}\Phi$. At the order $\frac{1}{N}$, these operators are responsible for the IR sub-contributions within the bubble chain when $k^{2}$ itself is in the hard region, as well as the IR contribution on the $(p-k)$ line when $p-k$ becomes soft (this means $k$ is hard). Clearly, the “tree-level” insertions within the bubble-chain are mutually independent and can be performed separately within the individual bubbles. Indeed, when $k$ is hard, only one line within each bubble can become soft, and the Taylor expansion for that soft momentum do not affect other bubbles. To perform the insertion responsible for the small $p-k$ region, one must Taylor expand the momentum flowing into the bubble-chain, but the condensates formed by making $p-k$ soft are still the tree-level condensates and do not correlate with the tree-level condensates formed within the bubble chain. At the external legs and any hard internal lines, there can also be an arbitrary number of tree-level $\frac{1}{N}\Phi^{2}$ insertions that always combine together with the $m_{r}^{2}$ insertion. All these tree-level condensates can be removed by simply setting $\mu^{2}=m^{2}$.
@@
[…]
@@
    -- $\frac{m^{2}}{(p-k)^{4}}=-\frac{m^{2}}{p^{2}}\sum_{n=0}^{\infty}\frac{k^{2n}}{(p^{2})^{n+1}}\frac{(n+1)\Gamma(1-\epsilon)\Gamma(\epsilon+n+2)}{\Gamma(2+\epsilon)\Gamma(-\epsilon+n+1)}\ ,$ (93)
    and with more mass insertions one can generalize without difficulty. In Appendix. B, we explain how such angular averages can be performed. These coefficient functions are clearly absent from any factorial growth. See Fig. 6 for a depiction of these operator contributions.
Figure 6: The contribution of the operator $O_{n}$ due to the soft $k^{2}$ region, in the presence of zero (upper) and one (lower) mass-insertions after expanding the $\frac{1}{(p-k)^{2}+m^{2}}$ at small $k$. The dashed line represents massless scalar propagator, black dot denotes the mass insertion and the gray-blob denotes the operator condensate $\langle O_{n}\rangle$. As the case of $\phi_{2}^{4}$, double-lines represent “amputated soft external fields” to keep track the operators associated to the coefficient functions.
Given the above and using the fact that all tree-level operators can be removed simply by setting $\mu=m$, the expansion at $\mu=m$ takes a very simple form
$S_{\rm NL}(p^{2})-\frac{g}{2(p^{2}+m^{2})^{2}}\langle:\Phi^{2}:\rangle=C_{{\rm I},{\rm NL}}(p^{2})+\sum_{n=1}^{\infty}\frac{\langle O_{n}\rangle}{(p^{2})^{n}}\frac{\,{}_{2}F_{1}\left(n,n;1;-\frac{m^{2}}{p^{2}}\right)}{(p^{2}+m^{2})^{2}}\ ,$ (94)
where $\langle O_{n}\rangle$ denotes the renormalized operator condensates at $\mu^{2}=m^{2}$. Notice that the disconnected contributions to $O_{1}$ also vanish at $\mu^{2}=m^{2}$ to the desired $\frac{1}{N}$ order. To obtain the last term in Eq. (94), one needs to expand the $\frac{1}{(p-k)^{2}+m^{2}}$ to the order $k^{2n-2}$ and perform the angular average, see Appendix. B for more details. For other values of $\mu$, there will be many tree-level operators as discussed above, but the coefficient function $C_{\rm I;{\rm NL}}$ remains essentially the same form with $m_{r}^{2}$ replacing $m^{2}$ in the pre-coefficients, while $\mu^{2}$ replacing $m^{2}$ in the logarithms. For coefficient functions of ${\cal O}_{n}$, one simply replacing $m$ with $m_{r}$. From the above, it is clear that the coefficient functions for $O_{n}$ contain no factorial growths at all. Thus, to search for factorial growths, it is sufficient to start from the coefficient function for the identity operator. We can write the $\frac{1}{N}$ order coefficient function of the identity operator for $S_{\rm NL}(p^{2})$ in the general form as
$C_{{\rm I},{\rm NL}}(p^{2})=\frac{1}{(p^{2}+m_{r}^{2})^{2}}\sum_{n=1}^{\infty}C^{(n)}_{\rm I,NL}(p^{2})\ ,$ (95)
$C^{(n)}_{\rm I,NL}(p^{2})=\frac{g^{n+1}}{(p^{2})^{n}}\sum_{k=0}^{n-1}\left(\frac{m_{r}^{2}}{g}\right)^{k}C_{n,k}(l_{\mu})\ ,$ (96)
@@ 3.1.1 Calculation of $C_{n,0}$: factorial growth from $\frac{1}{\epsilon}\times\epsilon$
Figure 7: The massless bubble chain diagram contributing to $C_{n,0}$ with $n=5$ without mass insertions (upper) and the IR subtractions (lower) excluding the one for $O_{n}$. Since the contraction of $O(N)$ indices is clear, we no longer separate the quartic interactions in the figure. The IR subtractions on multiple bubbles factorize and can be performed separately in each bubble. Similarly, the IR counter-term for the $p-k$ line also factorizes.
Given the above, let’s consider the calculation detail. For individual bubble chain, one needs the building block that already appeared earlier in Eq. (41)
$\frac{1}{2}C_{1}=\frac{\mu_{0}^{2-d}}{2}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}(p-k)^{2}}=\frac{1}{p^{2}}\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}f(\epsilon)\times\frac{-1}{4\pi\epsilon}\ ,$ (97)
@@
Given the above, the bare coefficient function (in the below we use hat to denote bare or un-subtracted coefficient functions) corresponding to the $n$-bubble chain diagram (depicted in the upper of Fig. 7 with $n=5$) reads
@@
The crucial property of the above is that it is analytic at $s=0$ and $\epsilon=0$. Now, the IR subtraction for each of the bubble can be performed by adding the operator UV poles $\frac{1}{4\pi\epsilon}$ in Eq. (19) for $\frac{1}{N}\Phi^{2}$ insertions replacing one of the two propagators in $C_{1}$ (clearly there are two ways, thus the subtraction is $2\times\frac{1}{2}\times\frac{1}{4\pi\epsilon}\times\frac{1}{p^{2}}$, where $\frac{1}{p^{2}}$ is another propagator Taylor-expanded to the zeroth order):
$\frac{1}{2}C_{1}\rightarrow\frac{1}{p^{2}}\times\frac{-1}{4\pi\epsilon}\left(\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}f(\epsilon)-1\right)\ .$ (102)
Therefore, the bubble-chain-partly-renormalized diagram can be obtained by replacing $s\rightarrow(n+1-k)\epsilon$ with $k$ ranging from 00 to $n$ and summing over $k$ using the binomial theorem:
$R_{b}\hat{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}\ .$ (103)
Notice we used the notation $R_{b}F$ to denote the bubble-chain-partly-renormalized quantity $F$, for which the IR sub-divergences within the bubble chain (not including the entire bubble-chain) have been subtracted. See Fig. 7 for a depiction of the bubble-chain diagram and IR subtractions for $C_{n,0}$ with $n=5$. Now, one needs the standard summation formulas Palanques-Mestre and Pascual (1984); Beneke and Braun (1994, 1995), as well as the new one (this is actually related to the second Stirling number $S(n+1,n)$)
@@
To show this, one writes
$\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k+x)^{n+1}$
$=x(n+1)\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n}+\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n+1}$
$=x(n+1)n!+n!\binom{n+1}{2}=\frac{n!(n+1)(n+2x)}{2}\ .$ (105)
This is due to the fact that $\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)^{n+1}$ is nothing but $n!S(n+1,n)$ where $S(n+1,n)$ is the Stirling number of the second kind counting the partition of $n+1$ objects into $n$ unlabeled sets. Given this summation formula, one obtains up to all constant terms,
$R_{b}\hat{C}_{n,0}$
@@
This term must be canceled by subtracting the IR contribution from the $p-k$ propagator. Since here the IR counter term is again the UV pole $\frac{1}{4\pi\epsilon}$ for the $\frac{1}{N}\Phi^{2}$ operator, one simply needs to multiply this UV pole with the IR-subtracted bubble chain at the momentum $k=p$. Therefore, one has
$T_{p-k}R_{b}\hat{C}_{n,0}=-\frac{1}{4\pi\epsilon}\times\frac{1}{(4\pi)^{n}\epsilon^{n}}\left(f(\epsilon)\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}-1\right)^{n}\ .$ (109)
Here, $T_{p-k}F$ denotes the quantity obtained from $F$ such that the $p-k$ line have been replaced by the appropriate IR counter-term. Expanding the bracket using the binomial theorem, after using the summation formulas one reaches
$T_{p-k}R_{b}\hat{C}_{n,0}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{1}{\epsilon}\frac{d^{n}}{ds^{n}}H(\epsilon,s)\bigg{|}_{s=0}+\frac{n+2}{2}\frac{d^{n+1}H(\epsilon,s)}{ds^{n+1}}\bigg{|}_{s=0}\bigg{)},$
@@
To further simplify, one needs the small $\epsilon$ epansion
$H(\epsilon,s)=\left(\frac{\mu^{2}}{p^{2}}\right)^{s}-\frac{\epsilon}{12}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\left(12\ln\left(\frac{\mu^{2}}{p^{2}}\right)+\pi^{2}s\right)+{\cal O}(\epsilon^{2})\ .$ (111)
@@
which indeed cancel with that for $R_{b}C_{n,0}$. Given the above, the partly renormalized $C_{n,0}$, including all IR subtractions except the one for the operator $O_{n}$ reads
$R_{p}\hat{C}_{n,0}=R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (113)
It remains to add the operator poles for $O_{n}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n-2}\Phi^{2}$. Clearly, since all the high-order derivatives in $s$ are finite in $R_{p}C_{n,0}$, the operator poles must cancel the remaining $p$-independent divergences in the first term of $R_{b}C_{n,0}$ after the equality in Eq. (3.1.1). Therefore, one writes fully renormalized $C_{n,0}$ as
$C_{n,0}=R_{p}\hat{C}_{n,0}-\frac{1}{(4\pi)^{n+1}}Z_{n}\left(\frac{1}{\epsilon}\right)\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\ ,$ (114)
where $-\frac{Z_{n}g^{n+1}}{(4\pi)^{n+1}}$ denotes operator UV poles for the $O_{n}\rightarrow{\rm I}$ mixing that are at the order $g^{n+1}$. See Fig. 8 for a depiction of the IR subtractions corresponding to UV poles of $O_{n}$. Clearly, since the dimension of $O_{n}$ is $4+2n-2=2(n+1)$, this term has the highest power in $g$ in the UV poles of $O_{n}$. We further separate the high-order derivatives in $s$ and constant terms in $C_{n,0}$:
$C_{n,0}=C_{n,0}^{A}+C_{n,0}^{B}\ ,$ (115)
where $C_{n,0}^{B}$ contains all the high-order derivatives in $s$ and is finite (its final form will be given in Eq. (127)), while
$C_{n,0}^{A}=-\frac{1}{(4\pi)^{n+1}}\bigg{(}\frac{(-1)^{n}}{(n+1)\epsilon^{n+2}}G(n,\epsilon,0)+Z_{n}\left(\frac{1}{\epsilon}\right)\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{)}\ .$ (116)
Now, for $n\in Z_{\geq 1}$ one can simplifies the above as
$C_{n,0}^{A}=-\frac{1}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)\Gamma(n+\epsilon)}{\Gamma(1+\epsilon)\Gamma(n-\epsilon)}\bigg{(}Z_{n}+\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ .$ (117)
@@
$Z_{n}=-\text{poles of}\ \bigg{(}\frac{1}{\epsilon^{n+1}}\frac{2(-1)^{n}\sin(\pi\epsilon)\Gamma(-2\epsilon)}{(n+1)\pi\Gamma^{2}(1-\epsilon)}\bigg{)}\ ,$ (118)
@@
Clearly, this part is free from any factorial growth since the function $\frac{\sin(\pi\epsilon)\Gamma(-2\epsilon)}{\Gamma^{2}(1-\epsilon)}$ is analytic in a neighborhood of $\epsilon=0$, thus the Taylor expansion coefficients can only growth exponentially.
Figure 8: The IR counter-terms due to UV poles of the operator $O_{n}$, in the presence of zero (upper) and one (lower) mass-insertions after expanding the $\frac{1}{(p-k)^{2}+m^{2}}$ at small $k$. The crossed circle denotes the $g^{n+1}$ order operator UV poles of the $O_{n}\rightarrow{\rm I}$ mixing. These contributions should be added back to $C_{n,0}$ and $C_{n+1,1}$ respectively.
@@
Notice that the first two terms arise due to $\frac{1}{\epsilon}\times\epsilon$ effects. For $n=1$, one obtains $C_{1,0}^{A}=0$, while
@@
$C_{n,0}\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{\mu^{2}}{p^{2}}\bigg{)}\times\left(1+\frac{\mu^{2}}{2^{n+1}p^{2}}+..\right)\ .$ (128)
Notice that here we introduced the following notation to distinguish alternating and non-alternating asymptotics in $n$: for alternating asymptotics of a quantity $F(n)$ we write $F(n)|_{n\rightarrow\infty,(-1)^{n}}$, while for non-alternating one we write $F(n)|_{n\rightarrow\infty,+1}$. It is interesting to notice that the leading factorial growth is again one more power higher. Here there are two possibilities. The first possibility is that they cancel within the same coefficient function for the identity operator but with mass-insertions, if $\mu^{2}$ is identified as the mass. We will rule out this possibility after calculating the diagrams with mass-insertions. The second possibility is that they cancel with the operator contributions for $O_{n+1}=\frac{g^{2}}{4N}\Phi^{2}(i\partial)^{2n}\Phi^{2}$ and $O_{n+2}$…., but with the same bubble-chain number, because this diagram contains power-divergences for these operators. This would be consistent with the interpretation that the non-alternating asymptotics in Eq. (128) can be regarded as an IR renormalon.
@@
$C_{n,0}\bigg{|}_{n\rightarrow\infty,(-1)^{n}}=\frac{(-1)^{n}}{(4\pi)^{n+1}}\times n!\times\bigg{(}\frac{p^{2}}{\mu^{2}}\bigg{)}\times\left(1+\frac{p^{2}}{2^{n+1}\mu^{2}}+...\right)\ .$ (129)
Later we will show that they cancel with factorial growths for operators $O_{n-1},O_{n-2}...$ but with the same numbers of bubbles.